OpenStudy (anonymous):

Select the equation in which the graph of the line has a negative slope, and the y-intercept equals 10. Answer 10x + 5y = –25 5x – 10y = –12 20x + 5y = 50 2x + 6y = 10

OpenStudy (anonymous):

I think this is it 20x + 5y = 50

OpenStudy (anonymous):

Since 20x does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 20x from both sides 5y=-20x+50 move all the terms not containing y to the right 5y=50-20x 5y/5=50/5 - 20x/5 simplify the left side by cancelling out common factors y=50/5-20x/5 simplify y=10-4x Using the y=mx+b formula, m=-4 and b=10